A pump is delivering 50m3/hr of water with a discharge pressure of 3.5kg/cm2.The water is drawn from a sump where water level is 5 meter below the pump centre line.The power drawn by the motor is 9.5kw at 90% motor efficiency.Find out the efficiency of the pump.
1
Expert's answer
2020-04-29T13:31:10-0400

Discharge pressure = "P_2=3.5g \times 10^4 N /m^2"


Suction pressure = "P_1= \rho\times g\times h= 1000\times9.8\times 5=4.9 \times10^4 N/m^2"

Pump delievering at the rate of "50 m^3 \/hr" , Power driven=9.5 kW, "\eta_m=90" %,


As we know that,


"\eta_p= \frac{(P_2-P_1)Q}{P\times \eta_m}"


"\eta_p= \frac{(34.3-4.9)50\times 10^4}{9.5\times.9\times3600}"

"\eta_p= 47" %